A7. A recent study shows that the annual cost of maintaining a building in Accra averages
200 Cedis with a variance of 260 Cedis. If a tax of 30% is introduced on all items
associated with the maintenance of building i.e. everything is made 30% more expensive.
Calculate the standard deviation of the annual cost of maintaining a building in Accra
1
Expert's answer
2020-05-19T17:58:09-0400

"\text {Let Y represents the annual cost of maintaining}\\
\text { a building in Accra, then, E(Y)=200, V(Y)=260,}\\
\text {Let R represents the annual cost of maintaining}\\
\text { a building in Accra after adding a tax of 30\%,}\\
R=1.30Y\\
V(R)=V(1.30Y)=1.30^{2}V(Y)\\
=1.30^{2}(260)=439.4,\\
\therefore \sigma_{R}=\sqrt{V(R)}=\sqrt{439.4}=20.962"